News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 3 March 2006 9.00am EST

During the last week Chandra completed the observing schedule as planned.

During this week the EPHIN 27 rail voltage returned to its nominal range following an extended period at low levels due to high operating temperatures resulting from target observing attitudes. The behavior was expected based on an increase in the allowed maximum temperature for the device, and no performance impacts were observed.

The schedule of targets for next week is shown below and includes a follow-up observation of the Target of Opportunity GRB 060218.

Radiation Belts                      Mar 5
M81*                    ACIS-S/HETG
Abell 1689              ACIS-I
N1 (2 obs)              ACIS-I       Mar 6
HD 164058               HRC-I        Mar 7
GRB 060218              ACIS-S
Radiation Belts
N1 (6 obs)              ACIS-I       Mar 8
Abell 1689              ACIS-I
NGC 6205                ACIS-S       Mar 9
RXJ 1131-1231           ACIS-S       Mar 10
HD 100453               ACIS-S
Radiation Belts
NGC 6205                ACIS-S       Mar 11
V373 Cep                ACIS-S
RX J0337-3457           ACIS-I
BD +303639              ACIS-S/LETG
SDSS J144842.45+04      ACIS-S       Mar 12
IRAS14348-1447          ACIS-S


All spacecraft subsystems continued to support nominal operations.

Return to Status Reports